OpenStudy (ineedhelpnow):

HOW do i solve square root five plus square root 20? simplifying terms before combining them.

OpenStudy (anonymous):

3 Sqrt5

OpenStudy (anonymous):

\[\sqrt{5} +\sqrt{20} = \sqrt{5} + 2 \times \sqrt{5} = 3 \times \sqrt{5}\]

OpenStudy (asnaseer):

\[\sqrt{5}+\sqrt{20}=\sqrt{5}+\sqrt{4*5}=\sqrt{5}+2\sqrt{5}=3\sqrt{5}\]

OpenStudy (ineedhelpnow):

where'd the two come from?

OpenStudy (amistre64):

\[\sqrt{20}=\sqrt{4}{5}=\sqrt{4}\sqrt{5}=2\sqrt{5}\]

OpenStudy (asnaseer):

\[\sqrt{4*5}=\sqrt{4}*\sqrt{5}=2\sqrt{5}\]

OpenStudy (amistre64):

thats the way it should look lol

OpenStudy (ineedhelpnow):

oooooh! thanks!!!